For the velocity potential function, ϕ = x ^ 2 - y ^ 2 find the velocity components at the point (4, 5)


1
Expert's answer
2022-02-14T14:33:22-0500

"\phi=x^2-y^2"

Point (4,5)

We know that

"u_x=\frac{d\phi}{dt}=2x"

"u_y=\frac{d\phi}{dt}=-2y"

"u_x=2\times2=4"

"u_y=-2\times5=-10"

Velocity


"v=\sqrt{u_x^2+u_y^2}=\sqrt{4^2+(-10)^2}=\sqrt{16+100}=\sqrt{116}"

"|v|=10.77"